Write Kp for the reaction 2KCIO3 (s) 2KCI (s) 302 (g) If the pressure of the oxygen over the reaction mixture is doubled, what happers to the equilibrium constant?
Answer #1

Since Kp = (PO2)^3 ---------(1)

If pressure of oxygen is double then

Kp' = (2* PO2)^3 = 2^3 * (PO2)^3 = 8 *(PO2)^3 ----(2)

From (1) & (2)

Equilibrium constant will increase by 8 times if pressure of oxygen doubles.